# From Features to Reference Points: Lightweight and Adaptive Fusion for Cooperative Autonomous Driving

### v1

@@ 1 Introduction
Cooperative autonomous driving (CAD) refers to the coordination and collaboration between multiple autonomous vehicles (AVs), to improve safety, efficiency, and overall performance in a shared driving environment [automation2020taxonomy]. Despite its potential, developing reliable CAD systems remains challenging due to the heterogeneity of onboard computing and sensing systems, as well as limitations in current wireless communication networks. For example, models developed for one type of vehicle may not perform well on others, posing challenges for feature-fusion-based approaches [head, thornton2025real], which have recently become popular. Furthermore, effective cooperation depends on fast and reliable data exchange between vehicles, however, constraints in communication bandwidth would lead to incomplete information sharing, ultimately degrading system performance and safety. To overcome these challenges, it is crucial to develop a robust cooperative frameworks capable of handling the heterogeneity of vehicle models and adapting to real-time changes in network conditions.
@@ 1.1 Limitations of Prior Work
Current solutions to CAD can be divided into three categories: high-level, intermediate-level, and low-level data sharing [han2023collaborative, opv2v]. High-level sharing (e.g., SAE J2735 [sae]) is bandwidth-efficient but lacks detail and depends heavily on each vehicle’s perception quality. Intermediate-level methods improve accuracy through feature-map fusion [fcooper, v2vnet], but require high bandwidth and are suffer from model heterogeneity. Low-level sharing of raw sensor data [cooper] provides richest information, but is prohibitively expensive for real-world deployment. These limitations motivate a new cooperative paradigm that bridges the gap between high-level and intermediate-level sharing.
@@ 1.2 Proposed Solution
The concept of reference points is widely used in various classic transformer based models, including Deformable DETR [detr], Efficient DETR [yao2021efficient], BEVFormer [bevformer], and the Deformable Attention Transformer (DAT) [xia2022vision]. Reference points in transformer architectures show how important they are for helping models understand spatial relationships and interpret complex scenes more effectively. We discovered that even when vehicles share only their reference points, autonomous driving performance improves significantly compared to non-cooperative approaches. Overall, the findings suggest a promising direction: both reference points and latent features play key roles in CAD and can be used individually or together to enhance overall perception performance.
@@ 2 Related Work
End-to-End Autonomous Driving. End-to-end autonomous driving (E2EAD) unifies perception, prediction, and planning within a single framework, reducing cascading errors between modules [uniad, genad, stp3]. Recent works leverage BEV representations and multi-task training to jointly address tracking, occupancy prediction, motion forecasting, and planning [bevformer, zhang2025bridging, bench2drive]. While these approaches focus on semantic and geometric consistency, they typically consider single-vehicle settings and overlook challenges in cross-agent communication and bandwidth. In this work, we extend E2EAD to a cooperative setting, proposing a communication-efficient framework for heterogeneous multi-vehicle collaboration.
Cooperative Perception. Cooperative perception enables vehicles to share complementary information to overcome occlusion and limited field-of-view issues [fcooper, cooper, opv2v]. Early fusion exchanges raw sensor data [v2vnet], achieving complete sensing but requiring high bandwidth. Late fusion reduces communication by sharing detection outputs [han2023collaborative, fadili2025late], but provides limited information and depends heavily on each vehicle’s perception. Intermediate feature fusion shares BEV features for joint detection [fcooper, cobevt, yazgan2024survey, hu2022where2comm, DiscoNet], improving occlusion handling yet still incurring substantial bandwidth. Recent works such as mmCooper [liu2025mmcooper], ERMVP [zhang2024ermvp], and CodeFilling [hu2024communication] enhance efficiency and robustness, but rely on high-dimensional, task-specific features with limited interpretability.
Query Fusion for Cooperative Perception. Recent studies explore query fusion for cooperative perception, replacing high-dimensional BEV feature sharing with query-level collaboration [univ2x, cooptrack]. Queries capture rich instance-level semantics, including category, dynamics, trajectory, and interactions, enabling end-to-end cooperative perception with lower bandwidth and better focus on important objects. However, queries remain high-dimensional continuous embeddings, posing challenges for real-time, large-scale deployment. Additionally, queries from heterogeneous models may carry inconsistent semantics, complicating cross-agent fusion. Finally, query fusion often requires additional end-to-end training, increasing system complexity and deployment cost.
@@ 3.1 Overview
As illustrated in Fig. 2, the proposed RefPtsFusion framework consists of three primary components: the ego vehicle pipeline, the cooperating vehicle pipeline, and the cross-agent RefPtsFusion module. On the ego vehicle side, we adopt the end-to-end autonomous driving model UniAD as the baseline [uniad]. After Deformable DETR [detr] processes the BEV features generated by BEVFormer [bevformer], it produces object-level information, including position, size, and velocity. These representations are then fused with the corresponding information shared from the cooperating vehicle through the cross-agent RefPtsFusion module. The fused outputs are subsequently propagated to downstream modules (e.g., tracking, motion forecasting, and path planning) to accomplish various perception tasks and ultimately generate the final driving trajectory. We assume that the cooperating vehicle is capable of generating a set of reference points representing the centers of detected objects. This assumption is reasonable, as object detection serves as a fundamental capability in most autonomous driving systems. If the cooperating vehicle additionally provides auxiliary information, e.g., object velocity or size, the RefPtsFusion module can achieve more effective fusion performance. In this work, we employ Deformable DETR on the cooperating vehicle as an example to validate the effectiveness of the proposed framework; however, the framework itself is model-agnostic and can be seamlessly integrated with any object detection networks adopted by cooperating vehicles.
@@ 3.2 Cross-Agent Fusion
The core of RefPtsFusion lies in the cross-agent geometric alignment and association of reference points extracted from both the ego and cooperating vehicles. The reference points received from the collaborating vehicle are first transformed into the ego vehicle’s coordinate frame, using existing alignment methods (e.g., BB-Alignment [song2024bb]), the details of which are beyond the scope of this work. Once aligned, the transformed reference points are matched to their ego-vehicle counterparts, and instance-level fusion is performed to yield a consistent and unified set of fused reference points. Together with the reference points coordinates, any associated attributes, e.g., velocity and size, will undergo similar transformations and fusion guided by their matched reference points.
@@
Given a sender–ego pair $(S,E)$, the sender’s reference points $\mathbf{p}_{S}^{n}$ are transformed into the ego coordinate system using the extrinsic calibration [nuscenes] between vehicles, represented by a transformation matrix $\mathbf{T}_{S\rightarrow E}\in\mathbb{R}^{4\times 4}$, defined as
@@
The fused reference points are central to maintaining spatial alignment and temporal consistency across agents. They serve as refined spatial anchors for the Tracker module, guiding query updates within both the perception decoder and the Temporal Aggregation Network (TAN) [motr]. Serving as precise spatial anchors, each reference point defines the query’s location in the current frame and can be propagated or updated according to ego motion or estimated object velocity. During temporal attention, they direct the model to focus on relevant regions in subsequent frames, effectively narrowing the search space and improving cross-frame association accuracy. By fusing reference points instead of raw feature embeddings, the technique achieves explicit geometric alignment across agents and over time, thereby enhancing tracking performance.
@@ 3.3 Velocity- and Size-Aware Enhancements
Velocity-Aware Fusion. In the Tracker module [motr], each query maintains an estimated velocity vector, which is used to propagate reference points across frames and constrain cross-frame associations. This allows the model to predict the expected location of each instance in the next frame and match it with new detections. Integrating velocity information into RefPtsFusion will enable the ego vehicle to preserve temporal continuity after fusion, ensuring more accurate and consistent tracking.
@@ 3.4 Heterogeneous Backbones
Handling model heterogeneity in cooperative autonomous driving remains a significant challenge. In real-world scenarios, vehicles may employ various pipelines built on entirely different backbones, ranging from CNN-based [center-based, pointpillars] to Transformer-based architectures [bevformer, true-detr]. These models differ not only in their feature representations but also in the structure and semantics of intermediate embeddings. Therefore, designing a cooperative framework that works across heterogeneous models is inherently difficult without access to architectural knowledge.
@@ 3.5 Homogeneous Backbone
As a general cooperative framework, RefPtsFusion can also be applied to homogeneous backbones. When collaborating vehicles use the same perception backbone, cooperative performance can be further enhanced by exchanging high-dimensional features. Unlike conventional query- or feature-based fusion methods [univ2x, cooptrack], which aggregate all queries or features, we explicitly select the Top-$K$ queries guided by reference points. After transforming coordinates into the ego frame, each query from the sender is associated with a confidence score $c_{S}^{n}$ produced by the detection head. A query $\mathbf{q}$ consists of a positional and a semantic part, and only the top-$K$ queries $\{\mathbf{q}_{S}^{(1)},\dots,\mathbf{q}_{S}^{(K)}\}$, ranked by their confidence scores, are transmitted by the sender:
@@ 4.1 Experimental Setting
Our experiments are conducted on the M3CAD [m3cad], a large-scale dataset designed to support multi-vehicle end-to-end cooperative autonomous driving research. It contains 204 sequences, covering 30,000 frames and over 267K annotated instances across diverse traffic scenarios and challenging weather conditions. Most importantly, it supports a broad range of tasks from object detection and tracking to mapping, motion forecasting, occupancy prediction, and planning, thereby facilitating holistic end-to-end evaluation of cooperative driving systems.
Evaluation Metrics. To provide a rigorous and comprehensive assessment, we evaluate performance across five functional modules. For tracking, AMOTA, AMOTP [bernardin2006multiple], and IDS are used to measure detection accuracy, localization precision, and identity consistency. Mapping performance is assessed using IoU-lane and IoU-road to capture topological and geometric fidelity of reconstructed BEV maps. Motion forecasting is evaluated with minADE, minFDE, and MR [liang2020pnpnet, luo2018fast, peri2022forecasting] to quantify trajectory accuracy and miss rate, while occupancy prediction uses IoU to assess spatial accuracy and completeness of predicted occupancy grids [liu2021swin, zhang2022beverse]. For planning, average L2 error evaluates trajectory smoothness, stability, and driving safety [stp3].
@@
No fusion [uniad] 0.254 0.495 0.455 94.3 53.8 0.312 0.336 0.003 78.8 65.4 0.401
F-Cooper [fcooper] 0.720 0.680 0.816 - - - - - - - -
M3CAD [m3cad] 0.812 0.490 0.916 95.7 56.7 0.251 0.262 0.001 86.2 73.3 0.234
UniV2X [univ2x] 0.697 0.601 0.835 95.6 55.6 0.349 0.363 0.001 80.5 63.8 0.221
@@ 4.2 Quantitative Comparison
We compare our method against several representative baselines: No Fusion, an ego vehicle model without cooperation; F-Cooper [fcooper], the first feature-level cooperative perception approach; M3CAD [m3cad], the most recent end-to-end feature-level fusion framework; and UniV2X [univ2x], a query-level method that exchanges high-dimensional queries across vehicles.
Table 1 shows that RefPtsFusion achieves competitive perception and planning performance by sharing only reference points among vehicles. Compared with feature- and query-based fusion methods, RefPtsFusion achieves competitive tracking accuracy (0.671 AMOTA and 0.758 Recall), only 0.037 and 0.092 lower than Query Fusion (0.697 AMOTA and 0.835 Recall). It also performs comparably on other tasks, e.g., motion forecasting and occupancy prediction, and offers slightly better results on mapping accuracy (96.0% IoU-Road, 58.3% IoU-Lane). This may be because, during the first-stage training, TrackFormer (in UniAD [uniad]) was optimized using fused reference points, which provided additional supervisory signals to adjust the parameters of BEVFormer. As a result, the generated BEV features became more informative and structured, offering richer spatial cues that enhanced MapFormer’s performance on the mapping task.

### v2

@@ 1 Introduction
Cooperative autonomous driving (CAD) refers to the coordination and collaboration between multiple autonomous vehicles (AVs), to improve safety, efficiency, and overall performance in a shared driving environment [1]. Despite its potential, developing reliable CAD systems remains challenging due to the heterogeneity of onboard computing and sensing systems, as well as limitations in current wireless communication networks. For example, models developed for one type of vehicle may not perform well on others, posing challenges for feature-fusion-based approaches [23, 25], which have recently become popular. Furthermore, effective cooperation depends on fast and reliable data exchange between vehicles, however, constraints in communication bandwidth would lead to incomplete information sharing, ultimately degrading system performance and safety. To overcome these challenges, it is crucial to develop a robust cooperative frameworks capable of handling the heterogeneity of vehicle models and adapting to real-time changes in network conditions.
@@ 1.1 Limitations of Prior Work
Current solutions to CAD can be divided into three categories: high-level, intermediate-level, and low-level data sharing [8, 29]. High-level sharing (e.g., SAE J2735 [21]) is bandwidth-efficient but lacks detail and depends heavily on each vehicle’s perception quality. Intermediate-level methods improve accuracy through feature-map fusion [5, 26], but require high bandwidth and are suffer from model heterogeneity. Low-level sharing of raw sensor data [6] provides richest information, but is prohibitively expensive for real-world deployment. These limitations motivate a new cooperative paradigm that bridges the gap between high-level and intermediate-level sharing.
@@ 1.2 Proposed Solution
The concept of reference points is widely used in various classic transformer based models, including Deformable DETR [41], Efficient DETR [30], BEVFormer [16], and the Deformable Attention Transformer (DAT) [27]. Reference points in transformer architectures show how important they are for helping models understand spatial relationships and interpret complex scenes more effectively. We discovered that even when vehicles share only their reference points, autonomous driving performance improves significantly compared to non-cooperative approaches. Overall, the findings suggest a promising direction: both reference points and latent features play key roles in CAD and can be used individually or together to enhance overall perception performance.
@@ 2 Related Work
End-to-End Autonomous Driving. End-to-end autonomous driving (E2EAD) unifies perception, prediction, and planning within a single framework, reducing cascading errors between modules [10, 38, 9]. Recent works leverage BEV representations and multi-task training to jointly address tracking, occupancy prediction, motion forecasting, and planning [16, 35, 13]. While these approaches focus on semantic and geometric consistency, they typically consider single-vehicle settings and overlook challenges in cross-agent communication and bandwidth. In this work, we extend E2EAD to a cooperative setting, proposing a communication-efficient framework for heterogeneous multi-vehicle collaboration.
Cooperative Perception. Cooperative perception enables vehicles to share complementary information to overcome occlusion and limited field-of-view issues [5, 6, 29]. Early fusion exchanges raw sensor data [26], achieving complete sensing but requiring high bandwidth. Late fusion reduces communication by sharing detection outputs [8, 7], but provides limited information and depends heavily on each vehicle’s perception. Intermediate feature fusion shares BEV features for joint detection [5, 28, 31, 11, 15], improving occlusion handling yet still incurring substantial bandwidth. Recent works such as mmCooper [18], ERMVP [36], and CodeFilling [12] enhance efficiency and robustness, but rely on high-dimensional, task-specific features with limited interpretability.
Query Fusion for Cooperative Perception. Recent studies explore query fusion for cooperative perception, replacing high-dimensional BEV feature sharing with query-level collaboration [33, 39]. Queries capture rich instance-level semantics, including category, dynamics, trajectory, and interactions, enabling end-to-end cooperative perception with lower bandwidth and better focus on important objects. However, queries remain high-dimensional continuous embeddings, posing challenges for real-time, large-scale deployment. Additionally, queries from heterogeneous models may carry inconsistent semantics, complicating cross-agent fusion. Finally, query fusion often requires additional end-to-end training, increasing system complexity and deployment cost.
@@ 3.1 Overview
As illustrated in Fig. 2, the proposed RefPtsFusion framework consists of three primary components: the ego vehicle pipeline, the cooperating vehicle pipeline, and the cross-agent RefPtsFusion module. On the ego vehicle side, we adopt the end-to-end autonomous driving model UniAD as the baseline [10]. After Deformable DETR [41] processes the BEV features generated by BEVFormer [16], it produces object-level information, including position, size, and velocity. These representations are then fused with the corresponding information shared from the cooperating vehicle through the cross-agent RefPtsFusion module. The fused outputs are subsequently propagated to downstream modules (e.g., tracking, motion forecasting, and path planning) to accomplish various perception tasks and ultimately generate the final driving trajectory. We assume that the cooperating vehicle is capable of generating a set of reference points representing the centers of detected objects. This assumption is reasonable, as object detection serves as a fundamental capability in most autonomous driving systems. If the cooperating vehicle additionally provides auxiliary information, e.g., object velocity or size, the RefPtsFusion module can achieve more effective fusion performance. In this work, we employ Deformable DETR on the cooperating vehicle as an example to validate the effectiveness of the proposed framework; however, the framework itself is model-agnostic and can be seamlessly integrated with any object detection networks adopted by cooperating vehicles.
@@ 3.2 Cross-Agent Fusion
The core of RefPtsFusion lies in the cross-agent geometric alignment and association of reference points extracted from both the ego and cooperating vehicles. The reference points received from the collaborating vehicle are first transformed into the ego vehicle’s coordinate frame, using existing alignment methods (e.g., BB-Alignment [24]), the details of which are beyond the scope of this work. Once aligned, the transformed reference points are matched to their ego-vehicle counterparts, and instance-level fusion is performed to yield a consistent and unified set of fused reference points. Together with the reference points coordinates, any associated attributes, e.g., velocity and size, will undergo similar transformations and fusion guided by their matched reference points.
@@
Given a sender–ego pair $(S,E)$, the sender’s reference points $\mathbf{p}_{S}^{n}$ are transformed into the ego coordinate system using the extrinsic calibration [3] between vehicles, represented by a transformation matrix $\mathbf{T}_{S\rightarrow E}\in\mathbb{R}^{4\times 4}$, defined as
@@
The fused reference points are central to maintaining spatial alignment and temporal consistency across agents. They serve as refined spatial anchors for the Tracker module, guiding query updates within both the perception decoder and the Temporal Aggregation Network (TAN) [34]. Serving as precise spatial anchors, each reference point defines the query’s location in the current frame and can be propagated or updated according to ego motion or estimated object velocity. During temporal attention, they direct the model to focus on relevant regions in subsequent frames, effectively narrowing the search space and improving cross-frame association accuracy. By fusing reference points instead of raw feature embeddings, the technique achieves explicit geometric alignment across agents and over time, thereby enhancing tracking performance.
@@ 3.3 Velocity- and Size-Aware Enhancements
Velocity-Aware Fusion. In the Tracker module [34], each query maintains an estimated velocity vector, which is used to propagate reference points across frames and constrain cross-frame associations. This allows the model to predict the expected location of each instance in the next frame and match it with new detections. Integrating velocity information into RefPtsFusion will enable the ego vehicle to preserve temporal continuity after fusion, ensuring more accurate and consistent tracking.
@@ 3.4 Heterogeneous Backbones
Handling model heterogeneity in cooperative autonomous driving remains a significant challenge. In real-world scenarios, vehicles may employ various pipelines built on entirely different backbones, ranging from CNN-based [32, 14] to Transformer-based architectures [16, 4]. These models differ not only in their feature representations but also in the structure and semantics of intermediate embeddings. Therefore, designing a cooperative framework that works across heterogeneous models is inherently difficult without access to architectural knowledge.
@@ 3.5 Homogeneous Backbone
As a general cooperative framework, RefPtsFusion can also be applied to homogeneous backbones. When collaborating vehicles use the same perception backbone, cooperative performance can be further enhanced by exchanging high-dimensional features. Unlike conventional query- or feature-based fusion methods [33, 39], which aggregate all queries or features, we explicitly select the Top-$K$ queries guided by reference points. After transforming coordinates into the ego frame, each query from the sender is associated with a confidence score $c_{S}^{n}$ produced by the detection head. A query $\mathbf{q}$ consists of a positional and a semantic part, and only the top-$K$ queries $\{\mathbf{q}_{S}^{(1)},\dots,\mathbf{q}_{S}^{(K)}\}$, ranked by their confidence scores, are transmitted by the sender:
@@ 4.1 Experimental Setting
Our experiments are conducted on the M3CAD [40], a large-scale dataset designed to support multi-vehicle end-to-end cooperative autonomous driving research. It contains 204 sequences, covering 30,000 frames and over 267K annotated instances across diverse traffic scenarios and challenging weather conditions. Most importantly, it supports a broad range of tasks from object detection and tracking to mapping, motion forecasting, occupancy prediction, and planning, thereby facilitating holistic end-to-end evaluation of cooperative driving systems.
Evaluation Metrics. To provide a rigorous and comprehensive assessment, we evaluate performance across five functional modules. For tracking, AMOTA, AMOTP [2], and IDS are used to measure detection accuracy, localization precision, and identity consistency. Mapping performance is assessed using IoU-lane and IoU-road to capture topological and geometric fidelity of reconstructed BEV maps. Motion forecasting is evaluated with minADE, minFDE, and MR [17, 20, 22] to quantify trajectory accuracy and miss rate, while occupancy prediction uses IoU to assess spatial accuracy and completeness of predicted occupancy grids [19, 37]. For planning, average L2 error evaluates trajectory smoothness, stability, and driving safety [9].
@@
No fusion [10] 0.254 0.495 0.455 94.3 53.8 0.312 0.336 0.003 78.8 65.4 0.401
F-Cooper [5] 0.720 0.680 0.816 - - - - - - - -
M3CAD [40] 0.812 0.490 0.916 95.7 56.7 0.251 0.262 0.001 86.2 73.3 0.234
UniV2X [33] 0.697 0.601 0.835 95.6 55.6 0.349 0.363 0.001 80.5 63.8 0.221
@@ 4.2 Quantitative Comparison
We compare our method against several representative baselines: No Fusion, an ego vehicle model without cooperation; F-Cooper [5], the first feature-level cooperative perception approach; M3CAD [40], the most recent end-to-end feature-level fusion framework; and UniV2X [33], a query-level method that exchanges high-dimensional queries across vehicles.
Table 1 shows that RefPtsFusion achieves competitive perception and planning performance by sharing only reference points among vehicles. Compared with feature- and query-based fusion methods, RefPtsFusion achieves competitive tracking accuracy (0.671 AMOTA and 0.758 Recall), only 0.037 and 0.092 lower than Query Fusion (0.697 AMOTA and 0.835 Recall). It also performs comparably on other tasks, e.g., motion forecasting and occupancy prediction, and offers slightly better results on mapping accuracy (96.0% IoU-Road, 58.3% IoU-Lane). This may be because, during the first-stage training, TrackFormer (in UniAD [10]) was optimized using fused reference points, which provided additional supervisory signals to adjust the parameters of BEVFormer. As a result, the generated BEV features became more informative and structured, offering richer spatial cues that enhanced MapFormer’s performance on the mapping task.